Alternate Exercise A Diane Manufacturing Company is considering investing $500,ooo in new equipment with an estimated useful life of 10 years and no salvage value. The equipment is expected to produce $320,000 in cash inflows and $20o,ooo in cash outflows annually. The company uses straight- line depreciation, and has a 30% tax rate. a. Determine the annual estimated net income and net cash inflow. b. Calculate the payback period c. Calculate the accounting rate of return. Problem E Merryll, Inc., is considering three different investments involving depreciable assets with no salvage value. The following data relate to these investments:
Answer #1
Facts Given :
Initial Investment (0 Year): $ 5,00,000
Useful Life : 10 Years
Salvage Value: Nil
Annual Cash Inflows: $ 3,20,000
Annual Cash Outflows: $ 2,00,000
Method of Depreciation: Straight Line Method
Depreciation : $50,000
(Cost of Equipement- Salvage Value)/ Useful Life
Tax Rate: 30%
a) Determine the Annual Estimated Net Income
Particulars Net Estimated Income (in $)
Annual Cash Inflows 3,20,000
Less:Annual Cash Outflows 2,00,000
Less: Depreciation               50,000
Income before Taxes               70,000
Less: Tax @ 30%               21,000
Annual Estimated Net Income               49,000
Determine the Net Cash Flow:
Particulars Net Cash Flow (in $)
Annual Cash Inflows 3,20,000
Less:Annual Cash Outflows 2,00,000
Less: Depreciation               50,000
Income before Taxes               70,000
Tax @ 30%               21,000
Annual Net Income               49,000
Add: Depreciation               50,000
Annual Net Cash Inflows               99,000
b. Calculate the Pay Back Period : Initial Investment / Annual Cash Inflows
Pay Back Period : 5,00,000/ 99,000 = 5.05
Pay Back Period = 5.05
c. Calculate the Accounting Rate of Return : Average Annual Profit After Tax / Average Investment
Average Annual Profit after Tax = Total Expected After Tax Profits / No of Years
Average Annual Profit after Tax = 49000*10/ 10 = 49000
Average Investment= (Initial Investment + Salvage Value)/2
Average Investment: ($5,00,000+0)/2= 2,50,000
Accounting Rate of Return = 49,000/2,50,000*100 = 19.60%
Accounting Rate of Return = 19.60%